OpenStudy (anonymous):

If the lines y = 3x − 2, y = nx + 11, and y = −px + 1 are parallel, then what is p + n? Answer 5 0 -3 -2

OpenStudy (anonymous):

If the lines are parallel, all the slopes of the lines are equal to each other, thus 3 = n = −p n= 3, p = -3. n+p = 3-3 = 0. Answer: 0